OpenStudy (anonymous):

NEED HELP! 1 over x minus 2 over x squared minus x

myininaya (myininaya):

\[\frac{1}{x}-\frac{2}{x^2-x} ?\]

myininaya (myininaya):

Directions are to combine fraction and simplify if possible?

OpenStudy (anonymous):

yes thats it!

myininaya (myininaya):

First factor x^2-x and let me know what you get

OpenStudy (anonymous):

(x+x)(x-x)?

myininaya (myininaya):

x-x=0 so (x+x)(0)=0 so x^2-x doesn't equal 0

myininaya (myininaya):

hint: \[\frac{x}{x}=1 \\ \frac{x}{x}(x^2-x) =x(\frac{x^2-x}{x})=x(\frac{x^2}{x}-\frac{x}{x})\]

myininaya (myininaya):

Basically I see that x^2 and x have a common factor x so x^2-x=x( what goes here )

OpenStudy (anonymous):

x?

myininaya (myininaya):

x(what)=x^2 x(what)=x

myininaya (myininaya):

for the first blank it is x that is x(x)=x^2 but x(what)=x?

OpenStudy (anonymous):

x(1)=x

myininaya (myininaya):

yes so you have \[x^2-x \\ x(x-1)\]

myininaya (myininaya):

since x(x)=x^2 and x(-1)=-x

myininaya (myininaya):

\[\frac{1}{x}-\frac{2}{x(x-1)}\] now the only way we will able to combine these fractions is if they have the same denominator

myininaya (myininaya):

what factor are we missing on the bottom for the first fraction

OpenStudy (anonymous):

x-1

myininaya (myininaya):

yes so we are going to multiply first fraction by (x-1)/(x-1) like so \[\frac{x-1}{x(x-1)}-\frac{2}{x(x-1)}\]

myininaya (myininaya):

so now we can combine the fractions since they have the same bottom part

myininaya (myininaya):

\[\frac{(x-1)-(2)}{x(x-1)}=\frac{x-1-2}{x(x-1)}\]

myininaya (myininaya):

you can finish this up by replacing -1-2 with?

OpenStudy (anonymous):

-3?

myininaya (myininaya):

yah

myininaya (myininaya):

\[\frac{x-3}{x(x-1)}\] your answer could look like this or they could have maybe remultiplied the bottom back to what it was

myininaya (myininaya):

that is x(x-1) was x^2-x initially

OpenStudy (anonymous):

i got it! thank you so much!

myininaya (myininaya):

np